Hello! I am in need of an explanation for the following question I was assigned for homework. I have spent over an hour working on it but none of my solutions were correct. It goes: My motor boat normally travels at 24 kilometers per hour in still water. One day I traveled 36 kilometers against a constant current in a river. If I had traveled with the current, I would have traveled 48 km in the same time. How fast was the current? The answer is stated in the back of the book as 3 3/7 but does not give any explanation on how to get the answer. Please provide an explanation! Thank you.
Answer by stanbon(75887)

My motor boat normally travels at 24 kilometers per hour in still water. One day I traveled 36 kilometers against a constant current in a river. If I had traveled with the current, I would have traveled 48 km in the same time.
---------------------
Against current DATA:
dist = 36 km ; rate = 24 - c km/hr ; time = d/r = 36/(24-c) hr
-------------
With current DATA:
dist = 48 km ; tate = 24+c km/hr ; time = d/r = 48/(24+c) hr
------------------------
How fast was the current?
Equation:
time = time
36/(24-c) = 48/(24+c)
3/(24-c) = 4/(24+c)
3(24+c) = 4(24-c)
72 + 3c = 96 - 4c
7c = 24
---
current speed = 24/7 = 3.43 km/hr
-----------
Cheers,
Stan H.
-------------